# references.bib

### v1

@@ Abstract
This work performs a critical examination of the (in)compatibility of AI models and E2EE applications. We explore this on two fronts: (1) the integration of AI “assistants” within E2EE applications, and (2) the use of E2EE data for training AI models. We analyze the potential security implications of each, and identify conflicts with the security guarantees of E2EE. Then, we analyze legal implications of integrating AI models in E2EE applications, given how AI integration can undermine the confidentiality that E2EE promises. Finally, we offer a list of detailed recommendations based on our technical and legal analyses, including: technical design choices that must be prioritized to uphold E2EE security; how service providers must accurately represent E2EE security; and best practices for the default behavior of AI features and for requesting user consent. We hope this paper catalyzes an informed conversation on the tensions that arise between the brisk deployment of AI and the security offered by E2EE, and guides the responsible development of new AI features.
@@ 1 Introduction
In an era where so many forms of communication have moved online, widespread E2EE has become a critical underpinning for the security of communications, providing critical protections for journalists, dissidents, and activists, as well as the everyday family and social communications of billions. As underscored by independent experts from the UN Office of the High Commissioner on Human Rights, strong encryption is essential to fundamental human liberties including free expression, the right to hold opinions, the right to privacy, and the right to assembly [ohchr]. Beyond individual privacy, the ubiquitous nature of E2EE offers systemic protections for particular at-risk users, preventing their activities from appearing exceptional: widespread deployment of E2EE has made privacy the norm. This normalization of secure communications, in turn, makes it harder to distinguish whether individuals are engaged in frivolous or sensitive activities, and deters the targeting of individuals merely for using privacy technologies, resulting in a critical “network effect” [WhittakerInterview] that creates the conditions necessary for journalists, human rights defenders, activists, and others to communicate securely with less fear of surveillance and repercussions.
@@
First, the way people interact with AI models is changing. While they originally served as standalone applications, AI models are now increasingly incorporated into other everyday applications, including messaging applications, in the form of AI “assistants.” Interacting with these assistants is often baked into the user experience by default, made readily available as part of the application client (e.g., within a messaging app). Such integration creates systemic data flows at scale between previously separate systems, and accordingly raises security and privacy considerations not limited to the realm of E2EE.
Second, high-quality training data is becoming scarce. This has created a race between model developers, with tech companies under increasing pressure to tap any potential source of human-written content [villalobosposition]. That which is publicly accessible online has already been harnessed, so privately held data is naturally the next resort, and indeed, companies have been quietly changing their terms of service to enable training AI models on more and more of the user data they hold [tan2024tos]—a practice about which the U.S. Federal Trade Commission has already raised concerns [ftc2024tos]. At the time of writing, E2EE messaging data on major existing platforms remains off limits for AI training, but companies may be incentivized to reconsider these policies in the near future. Apple has very recently launched AI assistant features that can process E2EE messaging content (see Section 2.5), and other companies may be considering similar options.
Third, recent trends in AI appear to have shifted business incentives. Over the last decade, E2EE became a central part of the business model of online intermediaries. Business incentives to protect proprietary and user data were aligned with human rights protections. For example, in 2016, Apple and Facebook were lauded for their public resistance to government requests for access to E2EE data, including their refusal to comply with government requests to build systemic weaknesses into their encryption systems which would allow law enforcement access to messaging data [benner2016apple]. However, in the last couple of years we have seen a shift towards prioritizing AI features as a necessary component of modern applications, and a corresponding search for sources of data to power these, sometimes even at the cost of existing product features and profits [nyt-search-furor, forbes-customer-experience, arstechnica-profits]. At the same time, interestingly, research and reporting shows that consumers themselves can be much more skeptical of AI-powered products [kikuchi2023growing, cicek2024adverse].
@@
1.1 Summary of contributions
We systematically analyze the above question from both a technological and a legal perspective, combining the expertise of an multidisciplinary research team across cryptography and security, AI and LLMs, and technology law.
We identify the key confidentiality and integrity properties provided by E2EE, for both individual and systemic privacy considerations, which might be adversely impacted by feeding E2EE content into AI models. This requires a novel conceptualization of key properties of E2EE, ranging from the strictly definitional, to realities of E2EE deployments in practice and their divergences from the strict definitions, to essential properties that only hold when E2EE is used at scale, to many important properties that are widely associated with E2EE but not technically within the scope of formal definitions.
We then examine a wide range of detailed technical configurations that could fall under the broad umbrella of “feeding E2EE content to AI models,” taking into consideration the state of the art in cryptography and privacy technologies, as well as the latest technical developments in AI. We distinguish the use of E2EE content for inference, for training, or for both. We assess the capacity of each technical configuration—that is, of each approach to integrating AI features into an initially E2EE system—to uphold the key guarantees of E2EE. We conclude that some configurations cannot uphold these guarantees, while some others can.
Next, we overview relevant areas of law, and provide a detailed analysis of the circumstances under which E2EE service providers are likely to be able to offer AI features which use E2EE content, consistent with their legal obligations under current US and EU law alongside a brief note on other jurisdictions. We highlight areas of legal uncertainty, and provide a detailed exposition of pending legal (or quasi-legal) processes where relevant. In addition, looking beyond the limits of existing law, we discuss relevant critiques of current law and diverse academic theories of consent and privacy (legal and otherwise).
Putting together all of the above, our analysis yields the following key recommendations summarized here and elaborated in Section 7. Key terms are in italics; we provide detailed definitions in Section 3.
@@ 1.1 Summary of contributions
      Prioritize endpoint-local processing whenever possible.
@@
        No third party can see or use any E2EE content [^2] Or any derivative of E2EE content (i.e., any non-trivial function of content). See Section 3.1 for full definitions. without breaking encryption, [^3] This may sound paradoxical; see § 4.2 for detailed discussion of technical configurations that may achieve this condition. and
@@
Our analysis indicates that there would be serious risks—security risks and legal risks—involved in diverging from these recommendations and thus undermining the established operation of the current E2EE ecosystem according to privacy-by-default practices. Perhaps the greatest risk is in eroding E2EE as a systemic guarantee—an assumed condition of any application or service that is held to be E2EE—and turning it instead into merely a bonus feature that is routinely compromised in exchange for convenience.
Currently E2EE offers essential protections for billions of users communicating confidentially on WhatsApp, iMessage, Signal, and other platforms daily. Moreover, the widespread adoption of E2EE today offers critical systemic protections (“network effects”) for users because of its everyday use for more and less sensitive communications alike—protections that a special-purpose platform just for top-secret communication would not be able to provide. These “network effects” currently protect privacy and freedom of expression for at-risk groups and everyday users around the world.
E2EE is a context in which confidentiality is the core guarantee to be held above all others—confidentiality is the essence of the product. Though E2EE may be relatively unusual in this respect, we believe our inquiry in this paper resonates with themes in the broader debate about privacy and trust in AI.
@@
In Section 2 (Background), we overview the properties of E2EE (§ 2.1), how AI assistants work (§ 2.2), how AI assistants can process E2EE data (§ 2.3), and some background on trusted hardware (§ 2.4). In Section 2.5 (Apple Case Study), we provide a case study of Apple’s recent deployment of Apple Intelligence, which can process data from E2EE environments: the only publicly announced scheme of this type, to date. Section 3 provides definitions of key terms that we rely on throughout the paper (§ 3.1), including a taxonomy from “strict E2EE” to “not E2EE” (§ 3.2).
In Section 4 (Technical Implications), we analyze and present a range of technical configurations in which AI assistants might process E2EE data, and associated security/confidentiality implications. In Section 6 (Legal Implications), we overview different areas of US and EU law and how they may constrain the use of E2EE data as input for AI models. In Section 7, we offer our detailed recommendations on how and when messaging data can be fed into AI models consistent with E2EE, taking into careful consideration both technical possiblities and limitations, and existing legal frameworks. Section 8 provides discussion and Section 9 concludes (and includes a detailed note on author contributions).
@@ 2 Background I: Encryption, AI, and Trusted Hardware
This section covers technical background, focusing on end-to-end encryption (Section 2.1), AI assistants (Section 2.2), how it is possible for AI assistants to process E2EE content (Section 2.3), and trusted hardware (Section 2.4). Lastly, Section 2.5 gives a brief overview of Apple’s recently deployed system in which E2EE content can be processed by AI assistants.
@@ 2.1 Background on End-to-End Encryption (E2EE)
End-to-end encryption (E2EE) is a standard of security for communication in which only the sender and an intended recipient can read communications between them. In particular, even the platform $P$ cannot read these communications, despite $P$ still acting as the intermediary for communication: that is, communications still go from $S$ to $P$ then $P$ to $R$.
@@
We conclude this section with some terminology and follow it with a slightly more detailed and technically accurate explanation of how E2EE works. [^4] A complete exposition of how E2EE works is beyond the scope of this paper.
@@
Message content that is unencrypted is described as plaintext. Normal, readable messages that you type into iMessage, Signal, or WhatsApp are plaintexts. Despite the name, plaintexts can be any type of content: audio files, images, spreadsheets, and any other digital information. Encryption provide a technique to convert (or scramble) plaintexts into ciphertexts, which are unreadable except to the designated recipient (and the sender who wrote it). [^5] The terms plaintext and ciphertext can each be used either as a noun or as an adjective.
@@ 2.1.1 The E2EE Confidentiality Guarantee
The core privacy guarantee of an E2EE system is called confidentiality. [^6] We will hereafter use the term “E2EE confidentiality” in this paper to stress that we are referring to this particular technical property, and not the common use of the word. Informally, E2EE confidentiality states that plaintext messages handled by an E2EE system can only be seen by the sender and intended recipients: no one else, including law enforcement or the service provider, can learn anything about the plaintext messages [e2ee-definition-rfc]. Importantly, this guarantee is stronger than saying that nobody else can learn the plaintext messages: E2EE confidentiality requires that nothing can be learned about the content of an E2EE message, including the message itself, but also any derivative of the message content. [^7] The formal mathematical definition of security of encryption states that no adversary can any function of the message content, except for the message length in bits. Thus, E2EE confidentiality means that anyone besides the sender and intended recipient must not be able to learn, for example, what letter of the alphabet a message begins with, or whether the message contains a given word.
@@
E2EE confidentiality refers to a specific technical definition of confidentiality for messages, based on encryption—E2EE confidentiality (or E2EE security) is a term of art with a precise meaning, and does not ensure perfect “confidentiality,” “privacy,” or “security” for every colloquial meaning of the terms. In general use, these terms are very broad and mean different things in different contexts. [^8] For example, E2EE technologies do not prevent your friend from choosing to disclose the sensitive information you shared with them to others, even though this might seem to “violate confidentiality” or “violate your privacy” in a broad sense. Indeed, no technology can prevent this. See also Section 5 for more discussion.
Consider the context of messaging. The E2EE confidentiality guarantee is both stricter than one might expect, in some senses, and less strict in other respects. If strictly enforced, it is incompatible with many features common in messaging apps.
Strictness of the E2EE confidentiality guarantee
A strict reading of the E2EE confidentiality guarantee would be incompatible common features such as link previewing, where an app displays a “preview” of a link that is sent by E2EE messaging. In this case, the link itself is E2EE message content, and in order to fetch and display the preview, a third party must process that link. Some E2EE applications only display link previews upon explicit user opt-in, but some have it on by default.
Permissiveness of the E2EE confidentiality guarantee
The E2EE confidentiality guarantee covers only encryption of message contents, as noted above. For any other auxiliary information related to messaging, such as who is messaging whom and at what times, E2EE does not impose any requirements or provide any guarantees. To give a more complex example, in 2021, the Government of India introduced legislation mandating that messaging platforms provide message “traceability,” or methods for determining which user first shared particular content [internet_society_traceability_nodate]. Which user first shared content is not the content itself, and this feature does not violate E2EE confidentiality.
Many academic definitions of E2EE messaging consider a simple model where the only activity that occurs is literal messaging—sending messages between senders and receivers and within group chats. In practice, E2EE messaging apps—like any modern messaging app—are more richly featured (e.g., including options to back up your data and see link previews). How such additional features process message contents, and whether such processing respects E2EE confidentiality, has been a subject of long debate. Some common features of E2EE applications enhance confidentiality despite not being part of the basic definition of E2EE; other common features of E2EE applications may limit confidentiality (such as link previewing). [^9] Other additional features include, for example: user reporting, user blocking, importing contacts, stories, message backups, message forwarding, blurring explicit images, autocorrect, accessibility features (such as text-to-voice), and hands-free integration or carplay. It is an instructive exercise to consider how each of these do or do not have implications for confidentiality and E2EE confidentiality (depending on how they are implemented). While outside the scope of our work, there is a delicate balance (which has been the source of much debate ) to be struck between the deployment of usability features that increase adoption of E2EE systems (and, thus, help enhance the systemic guarantees discussed in Section 2.1.2) on the one hand and strict E2EE confidentiality on the other hand.
Next, in Sections 2.1.22.1.3, we present a range of technical and non-technical properties often associated with E2EE applications in practice, but which are distinct from the basic E2EE confidentiality guarantee.
@@ 2.1.2 Systemic Properties of E2EE Systems
Even if an E2EE application provides strong technical privacy guarantees (including and beyond E2EE confidentiality), the strength of the privacy guarantee it offers will also depend on its usage in practice.
Messaging applications that are widely used for diverse purposes offer additional protection because their use is unremarkable, and hence, the use of the application in itself does not draw attention. In economic terms, this is a network effect where the (privacy-related) utility of the application for any given user depends on other users. For example, suppose that a particular E2EE messaging app $A$ happens to have been developed by a developer who is an outspoken member of a particular political community and has primarily advertised $A$ in the social circles and publications of that community. Accordingly, $A$ grows such that the userbase disproportionately represents members of this political community. Even if the application has no flaws whatsoever in providing E2EE confidentiality (and other properties), knowing that a particular person uses application $A$ can divulge their political views, if only by association.
This type of example motivates the intuition that to realize the full privacy potential that E2EE can provide, the use of the application should be commonly used and unremarkable in the relevant context. This “unremarkableness” is not a technical property: it is instead an emergent social and contextual property of the system’s use as a whole considered over the entire userbase (relevant to a context of use [^10] E.g, widespread use in country $X$ may not be helpful for a user in country $Y$.). Existing security and privacy literature has discussed similar network effects in privacy technologies not limited to encrypted messaging, and noted their critical importance to engineering effective privacy technologies [DBLP:conf/weis/DingledineM06].
@@
2.1.3 Auxiliary Security Properties of E2EE Systems
Besides the core guarantees mentioned so far in this section, there are additional (cryptographic) security properties that E2EE applications often provide in practice. While distinct from basic E2EE confidentiality, these properties have become very commonly associated with, or even expected of, modern E2EE applications. Notable examples include:
  •   $\bullet$
    Forward secrecy: An adversary who has compromised the communication channel (e.g., by stealing the victims’ secret session key) cannot recover messages that were transmitted before the session was compromised.
  •   $\bullet$
    Post-compromise security: An adversary that has compromised the communicate channel cannot recover messages that are transmitted after the protocol “heals” itself (e.g., after session keys are rotated).
  •   $\bullet$
    Deniability: No party can prove that a message was authored by any particular user of the system, even if the prover was one of the intended recipients of the message.
  •   $\bullet$
    Unlinkability: Knowledge of the fact that a user authored a particular message reveals no information about whether they authored other messages.

### v2

@@ Abstract
This work performs a critical examination of the (in)compatibility of AI models and E2EE applications. We explore this on two fronts: (1) the integration of AI assistants within E2EE applications, and (2) the use of E2EE data for training AI models. We analyze the potential security implications of each, and identify conflicts with the security guarantees of E2EE. Then, we analyze legal implications of integrating AI models in E2EE applications, given how AI integration can undermine the confidentiality that E2EE promises. Finally, we offer a list of detailed recommendations based on our technical and legal analyses, including: technical design choices that must be prioritized to uphold E2EE security; how service providers must accurately represent E2EE security; and best practices for the default behavior of AI features and for requesting user consent. We hope this paper catalyzes an informed conversation on the tensions that arise between the brisk deployment of AI and the security offered by E2EE, and guides the responsible development of new AI features.
Related Shorter Resources: Blog-post summaries of this paper by some of the authors are available at the NYU DeTaIL Lab Blog [nyu-detail-blog] and Tech Policy Press [tpp-blog]. For examples of others’ commentary on our work, see Matt Green’s perspective in his blog A Few Thoughts on Cryptographic Engineering [matt-blog], and Stephen Thomas’s succinct coverage on the NYU Center for Data Science blog [nyu-cds-blog].
@@ 1 Introduction
In an era where so many forms of communication have moved online, widespread E2EE has become an essential underpinning for the security of communications, providing critical protections for journalists, dissidents, and activists, as well as the everyday family and social communications of billions. As underscored by independent experts from the UN Office of the High Commissioner on Human Rights, strong encryption is essential to fundamental human liberties including free expression, the right to hold opinions, the right to privacy, and the right to assembly [ohchr]. Beyond individual privacy, the ubiquitous nature of E2EE offers systemic protections for particular at-risk users, preventing their activities from appearing exceptional: widespread deployment of E2EE has made privacy the norm. This normalization of secure communications, in turn, makes it harder to distinguish whether individuals are engaged in frivolous or sensitive activities, and deters the targeting of individuals merely for using privacy technologies, resulting in a critical “network effect” [WhittakerInterview] that creates the conditions necessary for journalists, human rights defenders, activists, and others to communicate securely with less fear of surveillance and repercussions—protections that a special-purpose platform just for top-secret communication would not be able to provide. These “network effects” currently protect privacy and freedom of expression for at-risk groups and everyday users around the world.
@@
First, the way people interact with AI models is changing. While they originally served as standalone applications, AI models are now increasingly incorporated into other everyday applications and throughout devices, including messaging applications, in the form of AI “assistants.” Interacting with these assistants is often baked into the user experience by default, made readily available as part of the application client (e.g., within a messaging app). Such integration creates new systemic data flows at scale between previously separate systems, and accordingly raises security and privacy considerations not limited to E2EE.
Second, high-quality training data is becoming scarce. This has created a race between model developers, with tech companies under increasing pressure to tap any potential source of human-written content [villalobosposition]. That which is publicly accessible online has already been harnessed, so privately held data is naturally the next resort, and indeed, companies have been quietly changing their terms of service to enable training AI models on more and more of the user data they hold [tan2024tos]—a practice about which the U.S. Federal Trade Commission has raised concerns [ftc2024tos]. At the time of writing, E2EE messaging data on major existing platforms remains off limits for AI training, but companies may be incentivized to reconsider these policies in the near future. Apple and Samsung have very recently launched integrated AI features that can process E2EE content on their devices (see Section 2.5), and other companies may be considering similar options.
Third, recent trends in AI appear to have shifted business incentives. Over the last decade, E2EE became a central part of the business model of online intermediaries. Business incentives to protect proprietary and user data were aligned with human rights protections. For example, in 2016, Apple and Facebook were lauded for their public resistance to government requests for access to E2EE data, including their refusal to comply with government requests to build systemic weaknesses into their encryption systems which would allow law enforcement access to messaging data [benner2016apple]. However, in the last couple of years we have seen a shift towards prioritizing AI features as a component of modern applications, and a corresponding search for sources of data to power these, sometimes even at the cost of existing product features and profits [nyt-search-furor, forbes-customer-experience, arstechnica-profits]. At the same time, interestingly, research and reporting shows that consumers themselves can be much more skeptical of AI-powered products [kikuchi2023growing, cicek2024adverse].
@@
E2EE is a context in which confidentiality is the core guarantee above all others—confidentiality is the essence of the product. Though E2EE may be relatively unusual in this respect, our inquiry in this paper resonates with themes in the broader debate about security, privacy, and trust in AI.
Summary of contributions
We systematically analyze the above question from both a technological and a legal perspective, combining the expertise of an multidisciplinary team across cryptography/security, AI/LLMs, and technology law. Our contributions and recommendations illustrate that technological solutions and legal frameworks must work in tandem in order to achieve effective and enforceable safeguards.
First, we identify the key confidentiality and integrity properties provided by E2EE, for both individual and systemic privacy, which might be adversely impacted by feeding E2EE content into AI models (§3). This requires a novel conceptualization of key properties of E2EE, ranging from the strictly definitional, to realities of E2EE deployments and their divergences from the strict definitions, to essential properties that only hold when E2EE is used at scale, to auxiliary properties that are widely associated with E2EE.
Second, we examine a range of technical configurations for AI processing of E2EE data, taking into consideration the state of the art in cryptography and privacy technologies, as well as the latest technical developments in AI (§4). We assess the capacity of each such technical configuration to uphold the guarantees of E2EE. We conclude that some configurations cannot uphold these guarantees, while some others can.
Third, we overview relevant areas of law, and provide an analysis of the circumstances under which E2EE service providers are likely to be able to offer AI features which use E2EE content, consistent with their legal obligations under current US and EU law (§§56). We highlight areas of legal uncertainty, and discuss pending legal (or quasi-legal) processes.
Finally, our analysis yields the four key recommendations below (§7; italic key terms defined in §3.1). Our analysis indicates that there would be serious risks—security risks and legal risks—involved in diverging from these recommendations and thus undermining the established operation of the current E2EE ecosystem.
@@
      Prioritize endpoint-local processing where possible.
@@
        No third party can see or use any E2EE content [^2] Or any derivative of E2EE content (i.e., any non-trivial function of content). See Section 3.1 for full definitions. without breaking encryption, [^3] This may sound paradoxical; see Section 4.2 for detailed discussion of technical configurations that may achieve this condition. and
@@
In Section 2 (Background), we overview the properties of E2EE (§2.1), how AI assistants work (§2.2), how AI assistants can process E2EE data (§2.3), and some background on trusted hardware (§2.4). In Section 2.5, we discuss three examples of recently deployed AI assistants that can process data from E2EE environments. Section 3 provides definitions of key terms that we rely on throughout the paper (§3.1), including a taxonomy from “strict E2EE” to “not E2EE” (§3.2).
In Section 4 (Technical Implications), we analyze and present a range of technical configurations in which AI assistants might process E2EE data, and associated security/confidentiality implications. In Sections 56 (Legal Implications), we overview different areas of US and EU law and how they may constrain the use of E2EE data as input for AI models. In Section 7, we offer our detailed recommendations on how and when messaging data can be fed into AI models consistent with E2EE, taking into careful consideration both technical possiblities and limitations, and existing legal frameworks. Section 8 provides discussion and Section 9 concludes (and includes a detailed note on author contributions).
@@ 2 Background I: Encryption, AI, and Trusted Hardware
This section covers technical background: end-to-end encryption (Section 2.1), AI assistants (Section 2.2), how it is possible for AI assistants to process E2EE content (Section 2.3), and trusted hardware (Section 2.4). Section 2.5 gives a brief overview of three deployed systems in which AI assistants can process E2EE data: Apple Intelligence, Samsung’s new AI-integrated devices, and the integration of MetaAI with WhatsApp.
For readers knowledgeable about E2EE, TEEs, and AI, we suggest skipping to Sections 2.4.2 and 2.5.
@@ 2.1 Background on End-to-End Encryption (E2EE)
End-to-end encryption (E2EE) is a standard of security for communication in which only the sender and an intended recipient can read communications between them. In particular, even the platform $P$ cannot read these communications despite still acting as the intermediary for communication: that is, communications still go from $S$ to $P$ then $P$ to $R$.
@@
We conclude this section with some terminology and follow it with a slightly more detailed and technically accurate explanation of how E2EE works.
@@
Message content that is unencrypted is described as plaintext. Normal, readable messages that you type into iMessage, Signal, or WhatsApp are plaintexts. Despite the name, plaintexts can include any type of content: audio files, images, spreadsheets, and any other digital information. Encryption provide a technique to convert (or scramble) plaintexts into ciphertexts, which are unreadable except to the designated recipient (and the sender who wrote it). [^4] The terms plaintext and ciphertext can each be used either as a noun or as an adjective.
@@ 2.1.1 The E2EE Confidentiality Guarantee
The core privacy guarantee of an E2EE system is called confidentiality. [^5] We will hereafter use the term “E2EE confidentiality” in this paper to stress when we are referring to this particular technical property, and not the common use of the word. Informally, E2EE confidentiality states that plaintext messages handled by an E2EE system can only be seen by the sender and intended recipients: no one else, including law enforcement or the service provider, can learn anything about the plaintext messages [e2ee-definition-rfc]. Importantly, this guarantee is stronger than saying that nobody else can learn the plaintext messages: E2EE confidentiality requires that nothing can be learned about the content of an E2EE message, including the message itself, but also any derivative of the message content. [^6] The formal mathematical definition of security of encryption states that given the message length in bits, no computational adversary can guess any function of the message content with success non-negligibly better than random guessing. Thus, E2EE confidentiality means that anyone besides the sender and intended recipient must not be able to learn, for example, what letter of the alphabet a message begins with, or whether the message contains a given word. See [wired_e2ee_def, hale2022end, e2ee-definition-rfc, scheffler2023sokcontentmoderationendtoend] for more detailed definitions of end-to-end encryption.
@@
E2EE confidentiality (or E2EE security) refers to a specific technical definition of confidentiality for messages, based on encryption. It is a term of art with a precise meaning, and does not ensure perfect “confidentiality,” “privacy,” or “security” for every colloquial meaning of the terms. In general use, these terms are very broad and mean different things in different contexts. [^7] For example, E2EE technologies do not prevent your friend from choosing to disclose the sensitive information you shared with them to others, even though this might seem to “violate confidentiality” or “violate your privacy” in a broad sense. Indeed, no technology can prevent this. See also Section 5 for more discussion.
2.1.2 Auxiliary Security Properties of E2EE Systems
Besides the core guarantees mentioned so far in this section, there are additional (cryptographic) security properties that E2EE applications often provide in practice. While distinct from basic E2EE confidentiality, these properties have become very commonly associated with, or expected of, modern E2EE applications. Notable examples include:
  •   $\bullet$
    Integrity [unger2015sok]: Plaintext messages cannot be modified during transit in undetectable ways.
  •   $\bullet$
    Forward secrecy [green2015forward]: An adversary who has compromised the communication channel (e.g., by stealing the victims’ secret session key) cannot recover messages that were transmitted before the session was compromised.
  •   $\bullet$
    Post-compromise security [cohn2016post]: An adversary that has compromised the communicate channel cannot recover messages that are transmitted after the protocol “heals” itself (e.g., after session keys are rotated).
  •   $\bullet$
    Deniability [canetti1997deniable]: No party can prove that a message was authored by any particular user of the system, even if the prover was one of the intended recipients of the message.
  •   $\bullet$
    Unlinkability [unger2015sok]: Knowledge of the fact that a user authored a particular message reveals no information about whether they authored other messages.
@@ 2.1.3 Systemic Properties of E2EE Systems
Even if an E2EE application provides strong technical privacy guarantees (including and beyond E2EE confidentiality), the strength of the privacy guarantee it offers will also depend on its usage in practice. Messaging applications that are widely used offer additional protection because their use is unremarkable, and hence, the use of the application in itself does not draw attention. In economic terms, this is a network effect where the (privacy-related) utility of the application for any given user depends upon other users.
For example, suppose that a particular E2EE messaging app $A$ is known to have a userbase which disproportionately represents a particular political community. Even if $A$ provides perfect E2EE confidentiality, knowing that a particular person uses $A$ can divulge their political views, if only by association. This type of example motivates the intuition that to realize the full privacy potential that E2EE can provide, the application should be commonplace and its use should be unremarkable in the relevant context. This “unremarkableness” is not a technical property: it is instead an emergent social and contextual property of the system’s use as a whole considered over the entire userbase (relevant to a context of use [^8] E.g, widespread use in country $X$ may not be helpful for a user in country $Y$.). Existing security and privacy literature has discussed similar network effects in privacy technologies not limited to encrypted messaging, and noted their importance to engineering effective privacy technologies (e.g., Tor) [DBLP:conf/weis/DingledineM06]. This key privacy property of E2EE systems, which cannot be achieved by technical means alone, is attained by popular E2EE applications such as iMessage and WhatsApp today—namely, the network effect gained from wide usage such that usage in itself is unremarkable.